# Self-Organising Memristive Networks as Physical Learning Systems

### v1

@@ Abstract
Learning with physical systems is an emerging paradigm that seeks to harness the intrinsic nonlinear dynamics of physical substrates for learning. The impetus for a paradigm shift in how hardware is used for computational intelligence stems largely from the unsustainability of artificial neural network software implemented on conventional transistor-based hardware. This Perspective highlights one promising approach using physical networks comprised of resistive memory nanoscale components with dynamically reconfigurable, self-organising electrical circuitry. Experimental advances have revealed the non-trivial interactions within these Self-Organising Memristive Networks (SOMNs), offering insights into their collective nonlinear and adaptive dynamics, and how these properties can be harnessed for learning using different hardware implementations. Theoretical approaches, including mean-field theory, graph theory, and concepts from disordered systems, reveal deeper insights into the dynamics of SOMNs, especially during transitions between different conductance states where criticality and other dynamical phase transitions emerge in both experiments and models. Furthermore, parallels between adaptive dynamics in SOMNs and plasticity in biological neuronal networks suggest the potential for realising energy-efficient, brain-like continual learning. SOMNs thus offer a promising route toward embedded edge intelligence, unlocking real-time decision-making for autonomous systems, dynamic sensing, and personalised healthcare, by enabling embedded learning in resource-constrained environments. The overarching aim of this Perspective is to show how the convergence of nanotechnology, statistical physics, complex systems, and self-organising principles offers a unique opportunity to advance a new generation of physical intelligence technologies.
@@ 1 Introduction
Physical learning systems can be broadly described as physical substrates whose intrinsic nonlinear dynamics can be exploited for learning [1, 2]. Examples include analog electrical circuits, photonic and opto–electronic systems, as well as stochastic spintronic systems [3]. A particularly promising approach aims to harness native brain–like, or neuromorphic, properties exhibited by certain nanoscale electronic devices, such as memristors (see Box 1), with physical memory and neuromorphic dynamics [4]. Among the most promising examples of such systems are Self-Organising Memristive Networks (SOMNs). These physical networks are described as self–organising because under external electrical stimulation, they can spontaneously acquire specific novel structural and functional properties [5], effectively making them dynamically reconfigurable electrical circuits. These macroscale, observable properties emerge from the collective dynamical interactions between their constituent nanoscale elements, in the spirit of “More is Different” [6]. As highlighted in this Perspective, SOMNs are unique physical learning substrates because their ability to encode, store, and process information arises directly from their material and structural properties, positioning them at the intersection of condensed matter physics, neuromorphic computing, nanotechnology, and network science.
The importance of the structure and connectivity of biological neuronal networks to learning was recognised by some of the earliest work on artificial intelligence [7, 8, 9]. Rosenblatt’s perceptron introduced probabilistic learning and statistical separability as central to cognition; he envisioned the perceptron as a biologically–inspired model for emergent dynamics from random but structured physical wiring, while also noting practical limitations in scaling such architectures [8, 9, 10, 11]. These early insights motivated the first experimental realisations of SOMNs as scalable physical networks with neuromorphic architectures and dynamics achieved using disordered nano-electronic materials [12, 13, 14, 15, 16].
An active area of research in neuromorphic computing is the exploration of physical substrates that can be engineered to mimic the dynamics and efficiency of biological neurons and synapses [17]. In particular, memristor devices connected in a cross-bar array architecture map directly onto the bipartite graph structure of artificial neural networks (ANNs) and can thus serve as a physical instantiation of matrix–vector multiplications for efficient inference of ANN models [18, 19].
@@
In this Perspective, we posit that SOMNs are unique physical learning substrates because they mimic the structure–function relation of the brain, which is the archetypal non–equilibrium learning system [22, 23, 24, 25, 26, 27, 28]. Typically comprised of nanowires [12, 13, 20] or nanoparticles [29, 15, 16], SOMNs are formed via bottom–up synthesis, which naturally produces a high density of interconnections, far surpassing the limits of conventional top-down lithographic fabrication [14]. Their electronic architecture is also inherently stochastic [30], just like biological neuronal networks, i.e. no two single brains are identical.
@@
Figure 1 summarises the key features of SOMNs as physical learning systems. In contrast to conventional hardware computing systems (and many unconventional ones), SOMNs can process information efficiently without the need for direct addressability of individual components [1], similarly to biological neural networks. SOMNs can operate as physical learning systems by applying relatively simple learning schemes that leverage their adaptive dynamical responses to input electrical signals. They can be viewed as complex physical systems, as they produce higher-order nonlinear dynamics, criticality and phase transitions in the order parameter. These non-trivial dynamics emerge from the interplay between nanoscale transport dynamics (which generate high degrees of freedom in their internal resistance states) and heterogeneous, recurrent network connectivity [24, 31].
Box 1: Memristive devices The ideal memristor is defined by the relationship between flux ($\Phi$) and charge ($q$): $d\Phi=M\,dq,$ (1) where $M$, the memristance, depends on $\Phi$ or $q$. If $M$ is constant, the device behaves as a linear resistor; if $M$ varies, the memristor exhibits nonlinear behaviour. Memristors are typically characterized by a "pinched hysteresis loop" in their current-voltage (II-$V$) graph, which intersects at zero voltage and current [32]. A general description of a memristive device is: $\frac{dx}{dt}=F(x,u),\quad y=H(x,u)u,$ (2) where $x$ represents internal (physical) states, $u$ is the input ($V$ or II), and $y$ the output. The function $H(x,u)$ corresponds to the memristance or memductance [33]. The concept dates back to the 19th century with studies on arc lamps and early radio components. However, in 1971, Leon Chua formalized the memristor as the fourth fundamental circuit element, defining it as a device that “remembers" the history of past inputs. From a practical perspective, it is useful to think of a memristive device as a resistance (e.g. satisfying Ohm’s law $RI=V$) with a time dependent $R(t)$, whose value depends on the history of $V(t)$ (or $I(t)$, depending on the type of device).
@@
This Perspective is intended for researchers across several disciplines. Physicists can view SOMNs as physical substrates to study emergent phenomena [6], like criticality by means of dynamical systems and statistical mechanics frameworks. Material scientists and chemists may be motivated to explore new types of functional nanomaterials with similar properties (e.g. dynamic reconfigurability, long-range spatio–temporal correlations, as well as resistance to perturbations and self–healing, all key properties of adaptive, complex systems [34]). Computational neuroscientists may find value in SOMNs as an abiotic platform for gaining new insights into how coupled, physics–constrained synaptic dynamics produce multi–scale brain-like dynamics and process information [35, 36, 37, 38]. Computer scientists may instead view these physical systems as a potential platform for efficient machine learning applications and for exploring new approaches to solving computing problems.
@@
Beyond scientific interest, SOMNs address practical challenges in embedded edge intelligence [39]. Their ability to learn in situ under resource constraints makes them ideal for edge applications like dynamic sensing, robotics and autonomous systems, as well as continuous real-time monitoring and prediction (see Fig. 1). Indeed, in these edge applications, SOMNs offer a viable solution to AI’s energy crisis [40]. By integrating memory and processing in a single system, SOMNs bypass the von Neumann bottleneck and enable efficient learning directly at the physical interface. To understand how SOMNs can bridge material properties and learning, we examine the physical principles that underpin their brain-like dynamics and highlight how these systems have been implemented to date to perform meaningful learning tasks.
@@
The remainder of this Perspective is organised as follows. sec. 2 discusses the neuromorphic properties and dynamics of SOMNs, including the underlying transport physics that produces physical plasticity (sec. 2.2) and criticality (sec. 2.3). Sec. 3 presents a theoretical explication of SOMNs as complex physical systems, including theoretical methods for elucidating how memristive nonlinearity and memory combined with the physical constraints of Kirchhoff’s circuit laws induce highly non-trivial dynamics in the conductance (sec. 3.1), and how a mean-field model can be constructed (sec. 3.2). Sec. 4 discusses two main approaches to physical learning demonstrated with SOMNs: physical reservoir computing (sec. 4.1) and associative learning (sec. 4.2). We conclude with an outlook in sec. 5.
@@ 2 Brain-like dynamics in SOMNs
To understand how SOMNs may be used as physical learning systems, it is first necessary to examine the physical mechanisms behind their nonlinear dynamics.
Nanoscale transport dynamics and a heterogeneous network connectivity underpin the ability of SOMNs to produce emergent brain-like properties, including: plasticity – the ability to adapt to changing input signals; and avalanche criticality – signal propagation exhibiting long-range spatio-temporal correlations [24, 28]. This section describes how these dynamics arise in the context of the most well studied SOMNs, comprised of nanowire (NW) and nanoparticle (NP) networks. Both NW and NP networks show similar brain-like dynamics related to collective memristive switching effects (see Fig. 2). However, the underlying nanoscale transport physics in single NW-NW and NP-NP junctions is notably different.
@@ 2.1 Physics of ionic/electronic transport
The dynamics of NW networks are governed by ionic transport in the junctions between NWs. Here, NWs comprised of an electromigrating metal and a dielectric are assembled in a random arrangement where each contact region between two NWs forms a nanoscale electrical junction [41, 42]. Most commonly, NW networks employ silver as the electromigrating metal with either a polymer coating (e.g., Polyvinylpyrrolidone, PVP) [43, 24, 44] and/or metal chalcogenide (e.g., Ag2S [12, 45]), though other materials have also been explored (e.g. Ag-TiO2 [46], AgI [47], Ag2Se [48, 49]). In the nanoscale junctions, the electric field is high enough to drive the dissolution of atomic Ag to Ag+ that then electromigrates towards the opposite wire, where it nucleates back to metallic Ag, forming conductive nanofilaments in the junction [44]. Counter to this additive process, minimisation of the interfacial energy and local electromotive force effects (nanobattery effect) act to dissolve the nanofilaments, becoming the dominant process when the applied voltage is removed. Therefore, similarly to biological neurons where the synaptic strength is associated with changes in ionic flow and to membrane depolarisation and spontaneous repolarisation, NW junctions modulate their conductance depending on input history and are characterised by a short-term memory ranging from milliseconds to seconds. This mechanism is one of the most common memristive behaviours, at the basis of Electrochemical Metallization (ECM) memories and atomic switch devices [50]. In addition to the external electrical bias, the main intrinsic factors are ion mobility and redox rates [51].
Zooming out at the macroscopic level, NW networks are characterised by metastable states induced by the collective interactions among many memristive junctions where nanofilaments are continuously reshaped due to the local redistributions of electrical potential and current. Under voltage sweep stimulation, NW networks exhibit a collective hysteresis response characteristic of memristive systems (cf. Fig. 2(a)).
@@
When large external biases are applied or local inhomogeneities are present, Joule heating can induce the local breakdown of single nanowires, creating gaps where metal ions electromigrate and form nanofilaments, similarly to NW junctions [52]. While the short-term dynamics of such new memristive elements is likely indistinguishable from those of a nanowire junction, the network topology, along with all the local redistributions of electrical potential and current, is irreversibly changed, thus giving to the network a long-term memory characteristic. With a terminology borrowed from the human connectome, the two mechanisms are soundly referred to as reweighting, when an existing synapse (junction) is modulated in conductance, and rewiring, when a new synapse (gap) is created [44].
The conductivity mechanisms in NP networks differ from those in NW networks. There is currently no evidence for electrochemical ionic effects, at least in the most studied system comprised of tin NPs [53, 16] that are typically studied in vacuum. Interestingly, Au discontinuous films  [54] have similar mechanisms, in both air and in vacuum. Electrochemical effects might be expected in, for example, networks of Ag NPs (where electrochemical properties are expected to be similar to that in NW networks) [55], but detailed studies of the mechanisms are still required.
Figure 2: Memristive behaviour of SOMNs. (a) Memristive behaviour of a NW network under voltage sweep stimulation in two-terminal configuration (see inset), characterized by an hysteretic loop in the I-V plane typical of memristive systems. Adapted from ref. [44]. (b) Short-term plasticity effects in a NW network, characterized by potentiation of the conductance state during voltage stimulation followed by spontaneous relaxation after the end of stimulation. Adapted from [56]. (c) Self–similar conductance changes related to memristive switching events in NP networks. Top panel: 2400 s of experimental data; lower panels: successive segments of the data with 10 and 100 greater temporal magnification and greater magnification on the vertical scale. Adapted from Ref. [16].
@@
NP networks are typically fabricated in vacuum thanks to nanoscluster production systems that deposit the nanoparticles on a solid substrate. There they form a 2D layer of connected groups of particles separated by nanogaps[57, 53]. With this method, the surface coverage can be precisely tuned to poise the system near the percolation threshold. This is vital as it results in scale-free network structures and consequently to scale-free dynamics [58]. Interestingly, it has recently been reported [59] that the architectures of 3D percolating networks are substantially the same as those of 2D systems.
Typical metallic materials employed are Sn[57, 53, 60, 16, 61], Bi[62], Au[63, 64, 54, 65], Ag [55, 66, 67, 68, 69, 70], and Cu [71], but recent work in which devices were fabricated using Mo NPs shows the diversity of materials that can be used [72]. The most recent work on Ag devices [69, 70] provides a possible connection to work on NWN devices and suggests that the evolution of Ag network morphology strongly affects memristive / switching behaviour.
@@
The memristive mechanism in NP networks is the following: under an external electrical bias, the strong electric field in the gaps between NPs causes surface atoms to move [73]. At low electric fields this force creates hillocks on the surface of NPs that reduce the gap size, thus modulating the tunneling current [74]. At high electric fields (or if the field is applied for a long time) the hillocks extend all the way across the gaps and form filaments [57]. Once the bias is removed, the hillocks/filaments relax back due to interfacial energy minimisation. The filament formation and breaking processes can be characterised as neuron-like integrate-and-fire mechanisms[75] and timescales are typically in the microseconds range [76]. Filament formation / breaking leads to well-defined changes in conductance i.e. switching events. When one gap switches it causes a redistribution of voltage and current across the network, leading to avalanches of events that resemble those observed in biological neural networks [77, 78] (cf. sec. 2.3).
The proof that the transport is localized into atomic-scale filaments is given by quantum conductance effects, where conductance states are observed in multiples of the fundamental quantum of conductance G0= 2e2/h[57]. Note that quantum conductance effects have also been observed in single NWs [52] and in NW networks [43], as a consequence of the nearly atomic scale of memristive conducting filaments [79]. Note that, as described below, the low field hillock and high field filament modes are useful for different kinds of computation [80, 74, 61, 81, 82, 83, 84].
@@
Thus, charge migration and atomic rearrangement in the nanoscale junctions of NW or NP are responsible for the memristive switching dynamics in SOMNs. These properties result in self-adaptation to changes in the input signals (synapse-like plasticity) and as well as critical behaviour (neuron-like avalanches). These topics are explored in more depth in the next two subsections.

### v2

@@ Abstract
Learning with physical systems is an emerging paradigm that seeks to harness the intrinsic nonlinear dynamics of physical substrates for learning. The impetus for a paradigm shift in how hardware is used for computational intelligence stems largely from the unsustainability of artificial neural network software implemented on conventional transistor-based hardware. This Perspective highlights one promising approach using physical networks comprised of resistive memory nanoscale components with dynamically reconfigurable, self-organising electrical circuitry. Experimental advances have revealed the non-trivial interactions within these Self-Organising Memristive Networks (SOMNs), offering insights into their collective nonlinear and adaptive dynamics, and how these properties can be harnessed for learning using different hardware implementations. Theoretical approaches, including mean-field theory, graph theory, and concepts from disordered systems, reveal deeper insights into the dynamics of SOMNs, especially during transitions between different conductance states where criticality and other dynamical phase transitions emerge in both experiments and models. Furthermore, parallels between adaptive dynamics in SOMNs and plasticity in biological neuronal networks suggest the potential for realising energy-efficient, brain-like continual learning. SOMNs thus offer a promising route toward embedded edge intelligence, unlocking real-time decision-making for autonomous systems, dynamic sensing, and personalised healthcare, by embedding continuous learning in resource-constrained environments. The overarching aim of this Perspective is to show how the convergence of nanotechnology, statistical physics, complex systems, and self-organising principles offers a unique opportunity to advance a new generation of physical intelligence technologies.
@@ 1 Introduction
Physical learning systems can be broadly described as systems comprised of physical substrates whose intrinsic nonlinear dynamics can be exploited for learning [1, 2]. Examples include analog electrical circuits, photonic and opto–electronic systems, as well as stochastic spintronic systems [3]. A particularly promising approach aims to harness native brain–like, or neuromorphic, properties exhibited by certain nanoscale electronic devices, such as memristors (see Box 1), with physical memory and neuromorphic dynamics [4]. Among the most promising examples of such systems are Self-Organising Memristive Networks (SOMNs). These physical networks are described as self–organising because under external electrical stimulation, they can spontaneously acquire specific novel structural and functional properties [5], effectively making them dynamically reconfigurable electrical circuits. These macroscale, observable properties emerge from the collective dynamical interactions between their constituent nanoscale elements, in the spirit of “More is Different” [6]. As highlighted in this Perspective, SOMNs are unique physical learning substrates because their ability to encode, store, and process information arises directly from their material and structural properties, positioning them at the intersection of statistical physics, neuromorphic computing, nanotechnology, and complex systems.
The importance of the structure and connectivity of biological neuronal networks for learning was already recognised by some of the earliest work on artificial intelligence [7, 8, 9]. Rosenblatt’s perceptron introduced probabilistic learning and statistical separability as central to cognition; he envisioned the perceptron as a biologically–inspired model for emergent dynamics from random but structured physical wiring, while also noting practical limitations in scaling such architectures [8, 9, 10, 11]. These early insights motivated the first experimental realisations of SOMNs as scalable physical networks with brain–like architectures and dynamics achieved using disordered nano-electronic materials [12, 13, 14, 15, 16].
@@
In neuromorphic and brain–inspired computing, a current focus of research is the exploration of materials and devices that can be engineered to mimic the dynamics and efficiency of biological neurons and synapses [21, 22]. In addition, memristor devices connected in a cross-bar array architecture map directly onto the bipartite graph structure of artificial neural networks (ANNs) and can thus serve as a physical instantiation of matrix–vector multiplications for efficient inference of ANN models [23, 24]. In this Perspective we posit that SOMNs are unique physical learning substrates because they mimic the structure–function relation of the brain, which is the archetypal non–equilibrium dynamical learning system [25, 26, 27, 28, 29, 30, 18]. Typically comprised of nanowires [12, 13, 19] (NWs) or nanoparticles [31, 15, 16] (NPs), SOMNs are formed via bottom–up synthesis, which naturally produces a high density of interconnections, far surpassing the limits of conventional top-down lithographic fabrication [14]. Their electronic architecture is also inherently stochastic [32], just like biological neuronal networks, i.e. no two single brains are identical.
@@
Figure 1 summarises the key features of SOMNs as multiterminal platforms for physical learning. SOMNs exhibit internal conductance states that adaptively respond to external electrical signals (Fig. 1 (a)), which may be delivered directly when embedded at the sensor edge for a wide range of applications (Fig. 1 (b)). Most importantly, the response is associated with internal dynamical states. Fig. 1c,d and e show, respectively, scanning electron microscopy (SEM) images revealing the complex topology of SOMNs based on electrochemically grown NWs, randomly dispersed NWs, and randomly deposited NPs. In contrast to conventional hardware computing systems (and many unconventional ones), SOMNs can process information efficiently without the need for direct addressability of individual components [1], similarly to biological neural networks. SOMNs can operate as physical learning substrates by applying relatively simple learning schemes that leverage their adaptive dynamical responses to input electrical signals, similar to biological plasticity and in contrast to global weight training used in ANNs. SOMNs can be viewed as complex physical systems, as they produce higher-order nonlinear dynamics, criticality and phase transitions in the order parameter. These non-trivial dynamics emerge from the interplay between nanoscale transport dynamics (which generate high degrees of freedom in their internal resistance states) and heterogeneous, recurrent network connectivity [27, 33].
@@
This Perspective is intended for researchers across several disciplines. Physicists can view SOMNs as physical substrates to study emergent phenomena [6], like criticality, by means of dynamical systems and statistical mechanics frameworks. Material scientists and chemists may be motivated to explore new types of functional nanomaterials with similar properties (e.g. dynamic reconfigurability, long-range spatio–temporal correlations, as well as resistance to perturbations and self–healing, all key properties of adaptive, complex systems [34]). Computational neuroscientists may find value in SOMNs as an abiotic platform for gaining new insights into how coupled, physics–constrained synaptic dynamics produce multi–scale brain-like dynamics and process information [35, 36, 37, 38]. Computer scientists may instead view these physical systems as a potential platform for efficient machine learning applications and for exploring new approaches to solving computing problems.
Box 1: Memristive devices The ideal memristor is defined by the relationship between flux ($\Phi$) and charge ($q$): $d\Phi=M\,dq,$ (1) where $M$, the memristance, depends on $\Phi$ or $q$. If $M$ is constant, the device behaves as a linear resistor; if $M$ varies, the memristor exhibits nonlinear behaviour. Memristors are typically characterised by a "pinched hysteresis loop" in their current-voltage (II-$V$) graph, which intersects at zero voltage and current [39]. A general description of a memristive device is: $\frac{dx}{dt}=F(x,u),\quad y=H(x,u)u,$ (2) where $x$ represents internal (physical) states, $u$ is the input ($V$ or $I$), and $y$ the output. The function $H(x,u)$ corresponds to the memristance or memductance [40]. The concept dates back to the 19th century with studies on arc lamps and early radio components. However, in 1971, Chua formalised the memristor as the fourth fundamental circuit element, defining it as a device that “remembers" the history of past inputs. From a practical perspective, it is useful to think of a memristive device as a resistance (e.g., satisfying Ohm’s law $RI=V$) with a time dependent $R(t)$, whose value depends on the history of $V(t)$ (or $I(t)$, depending on the type of device). In practice, memristive devices can be experimentally realised by exploiting the physical mechanism of resistive switching in nanoscale devices [41]. Key classes of resistive switching memory (ReRAM) devices include i) electrochemical metallisation memories (ECM), where metal cations form/dissolve conductive filaments, ii) valence-change memories (VCM) where anion/vacancy motion modulates interfacial barriers and filaments [42, 43], and iii) phase-change memories based on reversible amorphous-crystalline structural transitions [44]. While variability and stochastic effects typical of experimental devices can be limiting for memory applications, they can also be exploited as a computational resource in neuromorphic devices [45]. The extension of the two-terminal memristive concept to generic multiterminal memristive systems, as SOMNs, is discussed in Ref. [46]
@@
Beyond scientific interest, SOMNs have the potential to address practical challenges in embedded edge intelligence [47]. Their ability to learn in situ under resource constraints makes them ideal for applications in robotics and autonomous systems, as well as for continuous real-time dynamic sensing and predictive analytics (cf. Fig. 1 (b)). Indeed, in these edge applications, SOMNs offer a viable solution to AI’s energy crisis [48]. SOMNs not only bypass the von Neumann bottleneck, by integrating memory and processing in a single system, but also enable efficient learning directly at the physical interface with reduced software complexity. To understand how SOMNs can bridge material properties and learning, we examine the physical principles that underpin their brain-like dynamics and highlight how these systems have been implemented to date to perform meaningful learning tasks.
@@
The remainder of this Perspective is organised as follows. Sec. 2 discusses the brain–like dynamics of SOMNs, including the underlying transport physics that produces physical plasticity (sec. 2.2) and criticality (sec. 2.3). Sec. 3 presents a theoretical explication of SOMNs as complex physical systems, including theoretical methods for elucidating how memristive nonlinearity and memory combined with the physical constraints of Kirchhoff’s circuit laws induce highly non-trivial dynamics (sec. 3.1), and how a mean-field model can be constructed (sec. 3.2). Sec. 4 discusses two main approaches to physical learning demonstrated with SOMNs: physical reservoir computing (sec. 4.1) and associative learning (sec. 4.2). We conclude with an outlook in sec. 5.
@@ 2 Brain-like dynamics in SOMNs
SOMNs are multiterminal memristive systems (Fig. 2(a)) where brain-like dynamics arise from nanoscale transport dynamics and network connectivity. Emergent brain-like properties include plasticity, e.g. the ability to adapt to changing input signals; and avalanche criticality, e.g. signal propagation exhibiting long-range spatio-temporal correlations [27, 18]. Although the underlying nanoscale transport processes in NW and NP networks differ, both types of SOMNs show similar brain-like dynamics related to collective memristive switching effects.
@@ 2.1 Physics of ionic/electronic transport
NW networks. The dynamics of NW networks are governed by ionic transport in the junctions between NWs. The NWs are comprised of an electromigrating metal and a dielectric and assemble into a random arrangement where a nanoscale electrical junction forms at each contact region at the intersection between two NWs [17, 49]. Most commonly, NW networks employ silver as the electromigrating metal with either a polymer coating (e.g., Polyvinylpyrrolidone, PVP) [50, 27, 51] and/or metal chalcogenide (e.g., Ag2S [12, 52]), though other materials have also been explored (e.g. Ag-TiO2 [53], AgI [54], Ag2Se [55, 56]). Under externally applied electrical stimulation, the strong electric field in nanoscale junctions can force Ag+ cations to electromigrate towards the opposite wire, where they nucleate back to metallic Ag, and form conductive nanofilaments across the junction [51]. Evidence that the transport is localised into atomic-scale metallic filaments is given by quantum conductance effects, where conductance states are observed in multiples of the fundamental quantum of conductance $G\textsubscript{0}=2e\textsuperscript{2}/h$ [50, 57].
@@
Counter to this additive process, minimisation of the interfacial energy and local electromotive force effects (nanobattery effect) act to dissolve the nanofilaments [58], becoming the dominant process when the applied voltage is removed. Therefore, similarly to biological neurons where the synaptic strength is associated with changes in ionic flow and to membrane depolarisation and spontaneous repolarisation, NW junctions modulate their conductance depending on input history and are characterised by a short-term memory ranging from milliseconds to seconds. This mechanism is one of the most common memristive behaviours, at the basis of electrochemical metallisation (ECM) memories and atomic switch devices [59]. In addition to the external electrical bias, the main intrinsic factors are ion mobility and redox rates [60].
Figure 2: Memristive and brain–like dynamics of SOMNs. (a) Schematic of a multiterminal SOMN device. (b) Example of memristive behaviour of a NW network measured under multiple AC cycles in a two-terminal configuration, showing hysteresis loops in the I-V plane typical of memristive systems; line colours (from brown to red) depict cycle number; the inset shows a schematic of the electrochemical metallisation mechanism underlying memristive switching at each junction. Adapted from Ref. [51]. (c) Conductance matrix of a multiterminal NW-based SOMN device where each matrix element represents the effective resistance between a pair of terminals. Adapted from Ref. [61]. (d) Example of synaptic plasticity responses in multiterminal NW-based SOMNs where a spatiotemporal input (left panel, with multiterminal SOMN schematic as inset) drives changes in synaptic conductance (right panel); circled areas highlight changes related to direct stimulation of the corresponding conductance pathway, where temporally correlated inputs results in paired pulse facilitation (PPF). Adapted from Ref. [62].
@@
Zooming out at the macroscopic level, NW networks are characterised by metastable states induced by the collective interactions among many memristive junctions where nanofilaments are continuously reshaped due to the local redistributions of electric potential and current. Under AC voltage sweeps, NW networks exhibit a collective hysteresis response characteristic of memristive systems (cf. Fig. 2b). When large external biases are applied, or local inhomogeneities are present, Joule heating can induce local breakdown in single nanowires, creating gaps where metal ions electromigrate and form nanofilaments, as in NW junctions [57]. While the short-term dynamics of such new memristive elements are likely indistinguishable from those of NW junctions (from the electrical point of view), the network topology, along with all local redistributions of electrical potential and current, is irreversibly changed, thereby imparting long-term memory characteristics to the network. Borrowing terminology from connectomics [63, 64], resistive switching in NW junctions and the creation of new memristive connections across single NWs can be associated with two forms of plasticity: reweighting, where an existing synapse (junction) changes its conductance, and rewiring, where a new synapse is created by forming a previously absent connection across a gap [51]. The collective spatiotemporal dynamics of multiterminal SOMNs can be experimentally investigated by monitoring the evolution of the conductance matrix of the system where each matrix element is the effective resistance between two terminals of the SOMN (cf. Fig. 2c) and/or through voltage maps that provides information about the spatial distribution of the voltage drop when a stimulation is provided. [61]
@@
NP networks. NP networks are typically fabricated in vacuum using nanoscluster production systems that deposit the NPs on a solid substrate. There, they form a 2D layer of connected groups of particles separated by nanogaps[65, 66]. With this method, the surface coverage can be precisely tuned to poise the system near the percolation threshold. This is vital as it results in scale-free network structures and consequently scale-free dynamics [67]. Interestingly, it has been found that the architectures of 3D percolating networks are similar to those of 2D systems [68]. Typical metallic materials employed are Sn[65, 66, 69, 16, 70], Bi[71], Au[72, 73, 74, 75], Ag [76, 77, 78, 79, 80, 81], and Cu [82], but recent work in which devices were fabricated using Mo NPs shows the diversity of materials that can be used [83]. The most recent work on Ag devices [80, 81] provides a possible connection to work on NW network devices and suggests that the evolution of Ag network morphology strongly affects memristive switching behaviour.
The dominant microscopic conduction mechanisms in NP networks can differ from those in NW networks. In the most widely studied NP systems, notably Sn NP networks measured in vacuum, there is currently no evidence that electrochemical ionic transport plays a role [66, 16]; instead, conductance is generally consistent with field-driven atomic rearrangements in nanogaps that modulate tunnelling and can evolve into filamentary bridges under stronger driving. Au discontinuous films exhibit closely related field-driven nanogap/filament dynamics both in air and in vacuum [74]. These physical mechanisms are in contrast to electrochemical mechanisms that are responsible for the similar effects with NW-based SOMNs [76]. For example, quantum conductance effects have also been observed in NP-based SOMNs [65].
[…]
@@
The memristive mechanism in NP networks is described as follows: under an external electrical bias, the strong electric field in the gaps between NPs causes surface atoms to move [85]. At low electric fields, this force creates “hillocks" on the surface of NPs that reduce the gap size, thus modulating the tunneling current [86]. At high electric fields (or if the field is applied for a long time), the hillocks extend all the way across the gaps and form filaments [65]. Once the bias is removed, the hillocks relax back due to interfacial energy minimisation, while filaments break due to electromigration effects [65]. The filament formation and breaking processes can be characterised as neuron-like integrate-and-fire mechanisms[87] and timescales are typically in the microseconds range [84]. Filament formation / breaking leads to well-defined changes in conductance i.e., switching events. When one gap switches, it causes a redistribution of voltage and current across the network, leading to avalanches of events that resemble those observed in biological neural networks [88, 89] (cf. sec. 2.3).
These field-driven morphological changes provide a natural internal state for each junction, with bias-induced growth narrowing the tunnelling barrier and post-bias relaxation widening it again. The junction conductance therefore reflects a history-dependent effective gap size. Because the conductance of NP junctions depends exponentially on the instantaneous tunnelling barrier width (i.e. on the effective gap size), even a modest, history-dependent change in protrusion height produces a strong nonlinear potentiation during stimulation, followed by a gradual decay after stimulation. It has been shown that the low field hillock and high field filament modes are useful for different kinds of computation [90, 86, 70, 91, 92, 93, 94].
@@
Charge migration and/or atomic rearrangement in the nanoscale junctions between NWs or NPs are responsible for the memristive switching dynamics in SOMNs. Together with their network connectivity, these mechanisms result in self-adaptation to changes in the input signals (synapse-like plasticity) and critical dynamics (neuron-like avalanches), which are discussed in more detail below.